OpenStudy (anonymous):

You have a 4-card deck containing a king, a queen, a jack, and 10. You pick a card, keep it, and then pick another card. What is the probability that you pick exactly 1 face card? (Face cards are jacks, queens, and kings.)

OpenStudy (anonymous):

I think I got it it's 3/4

OpenStudy (anonymous):

its 1/2

OpenStudy (anonymous):

They're are 3 face cards out of 4. So the chances of pulling out a face card is 3/4

OpenStudy (anonymous):

1 face = (3/4) * (1/3) + (1/4) * (3/3) 1 face = (1/4) + (1/4) 1 face ) = (1/2)

OpenStudy (anonymous):

Well ok